# Performance Characterization of Containers in Edge Computing

### v1

@@ Abstract.
This paper presents an empirical evaluation of container-based virtualization on embedded operating systems commonly used in Internet of Things (IoT) deployments. Focusing on platforms like the Raspberry Pi, we investigate the feasibility and performance implications of deploying Docker containers in resource-constrained edge environments. Our study employs both microbenchmarks (CPU, memory, and network profiling) and macrobenchmarks (AI-driven inference, sensor I/O workloads) to capture a comprehensive view of system behavior. The analysis is conducted on a custom-built physical testbed comprising Raspberry Pi devices equipped with environmental sensors and camera modules, enabling real-time deployment and measurement of representative IoT workloads. Through comprehensive quantitative analysis across a diverse suite of IoT tasks and real-time application services, we identify key overheads introduced by containerization and characterize the challenges specific to embedded IoT contexts—such as limited hardware resources, cold-start delays, and suboptimal I/O handling. Performance metrics include CPU utilization, memory faults, cache misses, network throughput, and latency. Our findings highlight trade-offs between isolation and efficiency and provide insights for optimizing container configurations to meet the real-time and reliability requirements of edge computing applications.
Containers, Edge computing, Performance characterization
journalyear: 2025ccs: Computer systems organization Embedded softwareccs: Software and its engineering Software product lines
@@ 1. Background and Motivation
Industry 4.0 and IoT growth are driving complex workloads to resource-constrained edge devices (sub-1GHz CPUs, $\leq$ 1GB RAM, slow flash storage) that must reliably host services, process real-time sensor data, and maintain security with infrequent updates. Containerization addresses these challenges by providing lightweight isolation and portability, making Docker essential for edge platforms like Azure IoT Edge and AWS Greengrass (edge2, ). However, the performance impact of containers on ARM-based edge hardware running real-world IoT workloads remains poorly understood, particularly for sensor-driven applications with strict timing requirements with limited memory and processing power. Resource constraints demand of low-cost, system-on-chip platforms (like Raspberry Pi) require container solutions that simultaneously achieve two goals: (1) robust isolation and migration to ensure secure, portable execution across heterogeneous edge nodes, and (2) strict real-time performance guarantees that maintain low latency and minimal memory overhead—often with sub-millisecond timing requirements for industrial and medical applications. Current research has yet to systematically evaluate these tradeoffs for practical IoT workloads, leaving a crucial knowledge gap as containerization becomes the de facto standard for edge computing.
State of the Art and Limitations: Existing research on container performance in edge computing has primarily focused on GPU-accelerated devices like NVIDIA Jetson platforms for training deep learning models (edge3, ). Although tools like Sysbench are useful for evaluating basic system performance (edge6, ), they fail to capture the complexity of real-world IoT inference tasks, which involve sensor I/O, encryption, and hardware interactions. Many IoT workloads are tightly coupled with physical I/O, introducing additional overhead when transmitting data from the embedded system’s I/O interfaces to containerized applications. In contrast, evaluating containers in real-time, sensor-driven pipelines—like environmental sensing or camera input—reveals how containerization affects end-to-end inference latency (including the data transmission and processing latency), CPU usage, cache misses, and memory faults on system-on-board (SoC) devices. High cache misses indicate inefficient memory access patterns that slow down computation, while frequent memory faults (especially page faults) can stall execution or trigger kernel overhead—both of which degrade performance on memory-limited edge platforms. With comprehensive profiling of diverse IoT workloads across key system performance metrics, developers can identify optimal configuration and tuning parameters to meet the real-time requirements of edge computing platforms like the Raspberry Pi (RasPi).
@@ 3.0.2. Macrobenchmarks
Table 3. Macrobenchmarking Inference Latency (in seconds) for AI-Driven Facial Recognition Application
App Mode / Method Docker Host
Facial Recognition Face-KNN 900 320
OpenCV 10 13
@@
Table 4. Macrobenchmarking IoMT Workloads Latency (in seconds)
@@
Figure 1. Cache misses: LZW(left) AES(right)
@@
Table 5. Macrobenchmarking IoMT Memory Faults (i.e. page faults) for IoMT’s LZW and AES Workloads (Host vs. Docker)
@@ 5. Conclusion and Future Work
This paper presents a detailed performance evaluation of Dockerized applications on resource-constrained platforms like the Raspberry Pi. Our analysis demonstrates that, in most cases, Docker introduces minimal performance overhead compared to native execution. For compute-intensive workloads, performance can be optimized by tuning resource constraints such as CPU and memory limits. We emphasize the importance of understanding workload-specific requirements to configure containers effectively, especially in edge and IoT settings. Looking ahead, several directions merit further exploration. First, lightweight container orchestration frameworks tailored for embedded environments are needed to manage multiple workloads efficiently. This includes investigating solutions like Docker Swarm and Kubernetes variants optimized for the edge. Second, sharing runtime engines between containers with similar resource and execution profiles could reduce overhead and improve memory utilization. Third, integrating Docker-based workflows with CI/CD pipelines for embedded systems and Real-Time Operating Systems (RTOS) would enable faster and safer deployment cycles. Overall, we hope this work encourages further research into redesigning container infrastructure to better meet the demands of real-time, multi-tenant edge computing platforms.
@@
Figure 2. Microbenchmark comparison of LXD containers versus native host execution on Raspberry Pi 4: (a) File creation latency for 1MB files, (b) Storage throughput scalability, (c) End-to-end latency for memory-bound operations, and (d) End-to-end latency for CPU-intensive workloads.

### v2

@@ Abstract.
Edge computing addresses critical limitations of cloud computing—such as high latency and network congestion—by decentralizing processing from cloud to the edge. However, the need for software replication across heterogeneous edge devices introduces dependency and portability challenges, driving the adoption of containerization technologies like Docker. While containers offer lightweight isolation and deployment advantages, they introduce new bottlenecks in edge environments, including cold-start delays, memory constraints, network throughput variability, and inefficient I/O handling when interfacing with embedded peripherals. This paper presents an empirical evaluation of Docker containers on resource-constrained edge devices, using Raspberry Pi as a representative platform. We benchmark performance across diverse workloads, including microbenchmarks (CPU/memory/network profiling) and macrobenchmarks (AI inference, sensor I/O operations), to quantify the overheads of containerization in real-world edge scenarios. Our testbed comprises physical Raspberry Pi nodes integrated with environmental sensors and camera modules, enabling measurements of latency, memory faults, I/O throughput, and cold-start delays under varying loads. Key findings reveal trade-offs between container isolation and edge-specific resource limitations, with performance degradation observed in I/O-heavy and latency-sensitive tasks. We identify configuration optimizations to mitigate these issues, providing actionable insights for deploying containers in edge environments while meeting real-time and reliability requirements. This work advances the understanding of containerized edge computing by systematically evaluating its feasibility and pitfalls on low-power embedded systems.
Containers, Docker, Edge computing, Performance characterization
conference: ; ; journalyear: 2025copyright: acmcopyrightccs: Computer systems organization Embedded softwareccs: Software and its engineering Software product lines
@@ 1. Background and Motivation
Industry 4.0 and IoT growth are driving complex workloads to resource-constrained edge devices (sub-1GHz CPUs, $\leq$ 1GB RAM, slow flash storage) that must reliably host services, process real-time sensor data, and maintain security with infrequent updates. Containerization addresses these challenges by providing lightweight isolation and portability, making Docker essential for edge platforms like Azure IoT Edge and AWS Greengrass (edge2, ). However, the performance impact of containers on ARM-based edge hardware running real-world IoT workloads remains poorly understood, particularly for sensor-driven applications with strict timing requirements with limited memory and processing power. Resource constraints demand of low-cost, system-on-chip platforms (like Raspberry Pi) require container solutions that simultaneously achieve two goals: (1) robust isolation and migration to ensure secure, portable execution across heterogeneous edge nodes, and (2) strict real-time performance guarantees that maintain low latency and minimal memory overhead, often with millisecond timing requirements for industrial and medical applications. Current research has yet to systematically evaluate these tradeoffs for practical IoT workloads, leaving a crucial knowledge gap as containerization becomes the de facto standard for edge computing.
2. Related Works
Existing container optimization approaches for edge computing fall into four categories, each with their respective limitations: (1) GPU-focused training benchmarks (edge3, ) and basic system tools (edge6, ) that fail to capture real-time application inference performance where sensors and cameras must communicate directly with containers under strict latency constraints for low-cost, standardized system-on-board platforms like Raspberry Pi. Similarly, other related works along this direction (edge1, ; edge8, ; edge9, ; edge10, ) have only evaluated isolated aspects (e.g. CPU/memory overhead or network latency) of containers in edge computing but miss end-to-end pipeline performance (sensor → containerized processing → actuation). (2) Resource-sharing solutions like CNTR (CNTR, ) and Pocket (pocket, ) tackle container bloat and startup latency by sharing common dependencies via daemon processes, but introduce trade-offs in isolation and security. (3) Monitoring systems like Colibri (colibri, ) enables fine-grained monitoring of containerized workloads, achieving significant reductions in SLA violations, but does not address the fundamental challenges of physical IO devices-to-container communication latency. Similarly, specialized optimizations for computer vision workloads, such as those presented by Tellez et al. (face, ) show promising results for containerized face recognition tasks. However, these approaches assume ideal network conditions and do not consider the varied network configurations provided by containers (host network mode, bridge network mode) that significantly impact performance in actual edge deployments. Moreoever, they do not even account for the complex interplay complex interplay between camera data acquisition, container communication overhead, and the inference execution which are important factors that dominate end-to-end latency in real-world edge applications.(4) Recent architectural innovations for edge containerization demonstrate promising directions but lack empirical validation under realistic conditions. Existing edge containerization architectures like ECAIoT (journal, ) and DAG-accelerator (sec2, ) show promise but lack empirical validation under realistic constraints. Our container benchmarking can reveal critical network and I/O bottlenecks, enabling hardware-aware tuning of the DAG protocol. Testing these architectures on our edge testbed exposes failure modes and guides practical optimizations for real-world deployment. A critical limitation across existing work is the lack of comprehensive benchmarking for end-to-edge inference pipelines, particularly those involving real-time communication between hardware peripherals and containerized applications. This gap is especially pronounced for system-on-board platforms, where flash storage limitations, USB/GPIO bottlenecks, and thermal throttling can drastically impact containerized workloads. Many IoT workloads are tightly coupled with physical I/O, introducing additional overhead when transmitting data from the embedded system’s I/O interfaces to containerized applications. In contrast, evaluating containers in real-time, sensor-driven pipelines—like environmental sensing or camera streams reveals how containerization affects end-to-end inference latency (including the data transmission and processing latency), CPU usage, cache misses, and memory faults on system-on-board (SoC) devices. High cache misses indicate inefficient memory access patterns that slow down computation, while frequent memory faults (especially page faults) can stall execution or trigger kernel overhead—both of which degrade performance on memory-limited edge platforms. With comprehensive profiling of diverse IoT workloads across key system performance metrics, developers can identify optimal configuration and tuning parameters to meet the real-time requirements of edge computing platforms like the Raspberry Pi (RasPi).
@@ 4.0.2. Macrobenchmarks
Table 3. Macrobenchmarking inference latency (in Seconds) for AI-driven facial recognition application.
App Method / Network Mode Docker Host
Facial Recognition Face-KNN-Bridge 900 320
Face-KNN-Host 320 110
OpenCV-Bridge 10 13
OpenCV-Host 9 3
@@
Table 4. Macrobenchmarking end-to-end latency (in Seconds) for IoMT (paper13, ) workloads.
@@
Figure 1. Macrobenchmarking cache misses for IoMT LZW(left) and AES(right) workloads.
@@
Table 5. Macrobenchmarking memory faults (i.e. page faults) for IoMT’s LZW and AES workloads
@@ 6. Conclusion and Future Work
This paper presents a detailed performance evaluation of Dockerized applications on resource-constrained platforms like the Raspberry Pi. Our analysis demonstrates that, in most cases, Docker introduces minimal performance overhead compared to native execution. For compute-intensive workloads, performance can be optimized by tuning resource constraints such as CPU and memory limits. We emphasize the importance of understanding workload-specific requirements to configure containers effectively, especially in edge and IoT settings. Looking ahead, several directions merit further exploration. First, lightweight container orchestration frameworks tailored for embedded environments are needed to manage multiple workloads efficiently. This includes investigating solutions like Docker Swarm and Kubernetes variants optimized for the edge. Second, sharing runtime engines between containers with similar resource and execution profiles could reduce overhead and improve memory utilization. Third, integrating Docker-based workflows with CI/CD pipelines for embedded systems and Real-Time Operating Systems (RTOS) would enable faster and safer deployment cycles. Overall, we hope this work encourages further research into redesigning container infrastructure to better meet the demands of real-time, multi-tenant and heterogeneous edge computing platforms.
@@
Figure 2. Microbenchmark comparison of LXD containers versus native host execution on Raspberry Pi 4: (a) File creation latency for 1MB files, (b) I/O throughput scalability, (c) End-to-end latency for memory-bound operations, and (d) End-to-end latency for CPU-intensive workloads.